# Intelligent Routing for Sparse Demand Forecasting: A Comparative Evaluation of Selection Strategies

### v1

@@ Abstract.
conference: The 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’25), August 3–7, 2025, Toronto, ON, Canadajournalyear: 2025ccs: Computing methodologies Modeling methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
In this paper, we address the challenge of selecting the optimal forecasting model by developing and comparing three distinct meta-model routing approaches. These routers are designed to intelligently select the most suitable forecasting method for each individual demand series from a diverse model bank, which includes classical methods (e.g., Croston(Croston, 1972), ETS(Hyndman et al., 2002)), ML models (LightGBM(Ke et al., 2017)), and state-of-the-art DL approaches (DeepAR(Salinas et al., 2019), PatchTST(Nie et al., 2023)). Our investigated routing mechanisms include: (1) a rule-based system leveraging specific demand characteristics (variability and autocorrelation), (2) a machine learning router (LightGBM) trained on a comprehensive set of extracted time series features, and (3) a deep learning router (InceptionTime(Ismail Fawaz et al., 2020)) which operates directly on the raw time series data. We extensively evaluate these routing strategies using the Favorita(Kag, [n. d.]) dataset, demonstrating their effectiveness and potential for significant accuracy improvements over conventional single-model forecasting methods.
@@ 3.3. Data Overview
3.4. Model Bank Construction and Evaluation
To create a diverse set of forecasting options for our routing system, we established a model bank comprising seven distinct forecasting methods. These models span a range of complexities and approaches, ensuring a broad coverage of potential time series dynamics. The bank includes:
  • Classical methods: Exponential Smoothing (ETS), Croston’s method (Croston, 1972), a Naïve (last value) benchmark, and Moving Average.
  • Machine learning models: LightGBM (Ke et al., 2017).
  • Deep learning models: DeepAR (Salinas et al., 2019) and PatchTST (Nie et al., 2023). It is important to note that due to computational constraints, the deep learning models (DeepAR and PatchTST) were trained with a limited number of iterations (100).
The primary objective of the model bank is not to forecast directly but to serve as a pool of experts from which our routers can select. To facilitate this, we first determined the optimal model for each time seriesa priori within our training set (19,646 series). Each of the seven models was individually trained and evaluated on every time series. This evaluation was conducted for two separate forecasting horizons: 14 days and 30 days.
@@ 6. Discussion
The problems or shortcomings we found in our study open up many interesting new questions and areas that researchers can explore in the future. Expanding the model bank to include a wider array of forecasting techniques, particularly newer hybrid or DL models, and performing more extensive hyperparameter tuning and training, would be beneficial. Developing more sophisticated routing architectures, potentially incorporating attention mechanisms or graph-based methods (if a relationship between series can be established), could further enhance selection accuracy. Exploring cost-sensitive learning for the routers, where misclassifying to a particularly poor model incurs a higher penalty, could improve robustness. Furthermore, implementing explainable AI (XAI) techniques could provide valuable insights into why a router selects a specific model, increasing trust and understanding. Additionally, since the LightGBM and InceptionTime routers output class probabilities, exploring their use to create weighted average ensemble forecasts—multiplying each model’s forecast by its predicted probability—offers another promising direction beyond single-model selection. Finally, extending the framework to an online learning setting, where routers can adapt to evolving demand patterns over time, presents a significant opportunity for real-world supply chain applications.

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 1st Workshop on ”AI for Supply Chain: Today and Future” @ 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 1st Workshop on ”AI for Supply Chain: Today and Future” @ 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Modeling methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
In this paper, we address the challenge of selecting the optimal forecasting model by developing and comparing three distinct meta-model routing approaches. These routers are designed to intelligently select the most suitable forecasting method for each individual demand series from a diverse model bank. This model bank was strategically assembled to include models with complementary strengths, spanning classical methods tailored for specific patterns (e.g., Croston(Croston, 1972), ETS(Hyndman et al., 2002)), a robust ML model (LightGBM(Ke et al., 2017)), and state-of-the-art DL approaches capable of learning complex dependencies (DeepAR(Salinas et al., 2019), PatchTST(Nie et al., 2023)). The necessity of routing among these models stems from the well-documented observation that no single forecasting method consistently outperforms others across the wide variety of demand patterns found in real-world retail data (Makridakis and Hibon, 2000). Our investigated routing mechanisms include: (1) a rule-based system leveraging specific demand characteristics (variability and autocorrelation), (2) a machine learning router (LightGBM) trained on a comprehensive set of extracted time series features, and (3) a deep learning router (InceptionTime(Ismail Fawaz et al., 2020)) which operates directly on the raw time series data. We extensively evaluate these routing strategies using the Favorita(Kag, [n. d.]) dataset, demonstrating their effectiveness and potential for significant accuracy improvements over conventional single-model forecasting methods.
@@ 3.3. Data Overview
3.4. Model Bank Formulation and Evaluation
3.4.1. Rationale for Model Selection
A critical component of our framework is the model bank, which serves as a pool of candidate experts for the router. The selection of models was guided by the principle of **complementarity**, aiming to include a diverse set of methods, each with distinct theoretical underpinnings and strengths suited to different types of time series patterns observed in sparse demand data. This diversity is essential for the routing concept to be effective, as it ensures that for any given time series, a potentially suitable ’expert’ exists within the bank. Our goal was not to create an exhaustive list of all possible forecasting models, but rather to construct a representative and varied pool to rigorously test the efficacy of different routing strategies.
The model bank comprises seven forecasting methods, categorized as follows:
  • Classical Statistical Models: This group forms a foundation of well-understood and interpretable methods.
    • Exponential Smoothing (ETS): Chosen for its robustness and effectiveness in capturing underlying trend and seasonality, making it a strong candidate for more regular, ’smooth’ demand patterns (Hyndman and Athanasopoulos, 2018).
    • Croston’s Method: Included specifically for its theoretical design to handle intermittent demand. It separately models the time between demands and the demand size, making it a specialized expert for ’intermittent’ and ’lumpy’ series (Croston, 1972).
    • Naïve and Moving Average: These serve as essential, computationally inexpensive baselines. Their inclusion provides a ”simple but robust” option, which can often outperform more complex models on highly erratic or unpredictable series.
  • Machine Learning Model:
    • LightGBM: Represents the class of gradient boosting machines, which have proven highly effective in forecasting competitions (Ke et al., 2017). It was chosen for its ability to model complex, non-linear relationships between a rich set of engineered time series features (as described in Section 3.2) and the forecast target. It acts as a feature-based expert in the bank.
  • Deep Learning Models: This category includes modern, state-of-the-art architectures capable of learning representations directly from raw time series.
    • DeepAR: A probabilistic forecasting model based on recurrent neural networks (RNNs). It was selected for its ability to learn global models from large panels of related time series and handle temporal dependencies (Salinas et al., 2019).
    • PatchTST: A Transformer-based model that has recently demonstrated state-of-the-art performance on long-sequence forecasting tasks (Nie et al., 2023). Its inclusion tests whether cutting-edge attention mechanisms, which analyze data in ’patches’, provide a distinct advantage for certain series within our dataset.
While this selection represents a strong and diverse pool, we acknowledge that the efficacy of the routing system is inherently tied to the strength of its constituent models. For instance, the deep learning models (DeepAR and PatchTST) were trained with a limited number of iterations (100) due to computational constraints. While this may underestimate their full potential, it creates a realistic testbed for the router, whose task is to select the best model given the available trained instances. The primary objective of this study is to demonstrate the value of the routing mechanism itself, and this curated bank provides the necessary diversity to validate that concept.
3.4.2. Best Model Identification for Router Training
With the model bank established, the next step is to generate the ground-truth labels required for training our supervised learning routers (LightGBM and InceptionTime). The primary objective of the model bank at this stage is not to forecast directly, but to serve as a pool of experts from which we can determine the optimal model for each time series a priori within our training set (19,646 series).
To achieve this, each of the seven models was individually trained and evaluated on every time series. This evaluation was conducted for two separate forecasting horizons: 14 days and 30 days.
@@ 6. Discussion
The problems or shortcomings we found in our study open up many interesting new questions and areas that researchers can explore in the future. Expanding the model bank to include a wider array of more diverse or fully-tuned forecasting techniques, particularly newer hybrid or DL models, would be a valuable next step. Developing more sophisticated routing architectures, potentially incorporating attention mechanisms or graph-based methods (if a relationship between series can be established), could further enhance selection accuracy. Exploring cost-sensitive learning for the routers, where misclassifying to a particularly poor model incurs a higher penalty, could improve robustness. Furthermore, implementing explainable AI (XAI) techniques could provide valuable insights into why a router selects a specific model, increasing trust and understanding. Additionally, since the LightGBM and InceptionTime routers output class probabilities, exploring their use to create weighted average ensemble forecasts—multiplying each model’s forecast by its predicted probability—offers another promising direction beyond single-model selection. Finally, extending the framework to an online learning setting, where routers can adapt to evolving demand patterns over time, presents a significant opportunity for real-world supply chain applications.